# Toward Relative Positional Encoding in Spiking Transformers

### v1

@@ 4.2 XNOR-Based Spiking Self-Attention
In the standard self-attention mechanism of Transformer (Vaswani, 2017), the attention map is computed as the dot product between the query matrix $\mathbf{Q}$ and key matrix $\mathbf{K}$, i.e., $\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}$. The primary concept behind this approach is to use the dot product to measure the similarity between each token in $\mathbf{Q}$ and $\mathbf{K}$, which works effectively when both $\mathbf{Q}$ and $\mathbf{K}$ are continuous-valued matrices. However, when $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices, the dot product spike matrices, the dot product becomes inadequate for capturing the temporal or event-based interactions intrinsic to spike-based representations. As a result, the dot product operation in spiking self-attention, as outlined in Section 3.2, is not well-suited for spiking Transformers. To address this limitation, several researchers (Yao et al., 2023; Zhou et al., 2023a; Yao et al., 2024b) have proposed alternative attention mechanisms specifically designed for SNNs.
In this paper, we propose an XNOR-based spiking self-attention mechanism, which leverages the Hamming distance (See Section 3.4) to measure the similarity between each token in $\mathbf{Q}$ and $\mathbf{K}$. This approach enables the integration of the Gray-code-based relative positional encoding method into the attention mechanism, allowing for more effective capture of positional information in spike-based representations. Formally, we modify Equation 5 as follows:
$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K}),$ (11)
@@
where $\lnot$ denotes the Not operation, $\oplus$ denotes the XOR operation, and $D$ represents the channel dimension. It is important to note that every token in $\mathbf{Q}$ will perform XOR $\oplus$ with every token in $\mathbf{K}$, so we sum over the channel dimension $D$ to get $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, shown in Figure 2 (a). Compared to the dot product, the XNOR operation significantly increases the values in $\mathbf{AttnMap}\in\mathbb{N}_{0}$. Therefore, the scale factor $\tau$ in Equation 6 should be set to a smaller value or treated as a learnable parameter, ensuring that the firing rate of $\mathcal{SN}$ does not become excessively large. We will empirically demonstrate in the Experiments Section that this XNOR modification does not negatively impact the performance of the vanilla spiking self-attention.
4.3 Approximate RPE Method Based on Gray Code
We modify the SSA module by replacing the dot product operation with the XNOR operation, enabling the use of Hamming distance to measure the similarity between tokens. In this context, we propose that Gray Code can serve as an approximate approach to relative positional encoding for SNNs. This is supported by Theorem 4.1, which demonstrates the effectiveness of Gray Code in partially capturing positional relationships within the framework of XNOR-based spiking self-attention.
Theorem 4.1.
@@ 4.3 Approximate RPE Method Based on Gray Code
For two integers differing by $2^{n}$ in decimal, their Gray Code representations have a consistent Hamming distance of $1$ for $n=0$, and $2$ for $n\geq 1$, i.e., for any non-negative integer $a$ and $n$:
Proof.
Please refer to Appendix A for detailed proof. ∎
From Theorem 4.1, we can infer that Gray Code ensures the consistency of distance representations for every $2^{n}$ ($n\geq 0$) relative distance. This property makes Gray Code an ideal approximate RPE method for SNNs, as it not only partially captures relative positions but also preserves the binary nature (00 and $1$) nature of SNNs.
@@
In our implementation, we concatenate the Gray Code representations of each position index to both the query matrix $\mathbf{Q}\in\mathbb{N}_{0}^{T\times L\times D}$ and key matrix $\mathbf{K}\in\mathbb{N}_{0}^{T\times L\times D}$, leaving the remaining operations unchanged. Formally, the attention map $\mathbf{AttnMap}$ is computed as follows:
$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot([\mathbf{Q}\parallel G(\bm{l})]\oplus[\mathbf{K}\parallel G(\bm{l})]),$ (13)
@@
where $G(\cdot)$ represents the function that converts non-negative integers into their corresponding Gray Code representations. The vector $\bm{l}$ denotes an array of position indexes, specifically $[0,1,2,\dots,L-1]$, where $L$ is the sequence length of both $\mathbf{Q}$ and $\mathbf{K}$. $\parallel$ denotes concatenation on the channel dimension $D$. We name this method as Gray-PE.
4.4 Extended Variant on Attention Maps
@@
Although Gray-PE can partially capture relative distances, it faces significant challenges when the input sequence length is large or when the downstream task is highly sensitive to long-range dependencies. For instance, when the input sequence length reaches $10^{2}$, Gray-PE can only distinguish a limited range of relative distances. To address this issue, we propose an extension of Gray-PE, referred to as Log-PE. Specifically, we simulate Equation 7 by directly adding a pre-assigned relative position map, denoted as $\mathbf{R}_{ij}$, to the attention map produced by the XNOR-based SSA operation. To be formal:
$\small\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K})+\mathbf{R}_{i,j}$ (14)
where
@@
$\mathbf{R}_{i,j}=\begin{bmatrix}R_{i,j}\end{bmatrix}=\begin{bmatrix}\lceil\log_{2}(\frac{L-1}{|i-j|})\rceil\end{bmatrix}.$ (15)
@@
Here, $\lceil$.$\rceil$ . ⌉ denotes the round-up function, $L$ is the sequence length, and $i,j$ Figure 2 (c) shows an illustration of Log-PE, indicating that the farther relative distance, the smaller $R_{i,j}$. Since the original $\mathbf{AttnMap}\in\mathbb{N}_{0}$ is a matrix composed of non-negative integers, our primary goal is to ensure accurate relative distance consistency while preserving the effectiveness of spiking self-attention. If we set the $R_{i,j}$ as $\frac{L-1}{|i-j|}$, we could obtain a complete relative positional encoding for the XNOR-based SSA. However, we choose not to pursue this solution because, for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|}$ would severely distort the original representations in the spiking self-attention map. Therefore, using the logarithmic form $R_{i,j}$ represents a trade-off that balances the values of the attention map and relative positional encoding.
@@ 4.5 Two-Dimensional Form for Image Patches
CNN-based SNN models, such as Spiking VGG (Sengupta et al., 2019) and SEW-ResNet (Fang et al., 2021), do not incorporate the concept of “positional encoding” in their spike representations. In contrast, Vision Transformer (Dosovitskiy et al., 2021) reformulated traditional image classification into a patch-based approach, dividing images into smaller patches. Unlike 1D positional encoding, which only considers the linear sequence of patches, 2D RPE accounts for both the horizontal and vertical positions of the patches in the image grid. This ensures that the model can recognize the relative positions along a single axis and the crucial interactions between patches across both dimensions. In our implementation, we assign horizontal and vertical positions with independent dimensions to store the Gray Code. Formally, the attention map $\mathbf{AttnMap}$ is:
$\small\begin{split}\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(&[\mathbf{Q}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]\\ \oplus&[\mathbf{K}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]).\end{split}$ (16)
Here, $\mathbf{h}$ is the array of position indexes, specifically $\mathbf{h}=[0,1,2,\dots,h-1]$, where $h$ denotes the maximum patch index along the height axis. Similarly, $\mathbf{w}$ is along the width axis. We show our 2D form in Figure 2 (d).
Table 1: Experimental results of time-series forecasting on $4$ benchmarks with various prediction lengths $6,24,48,96$. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. “w/o” denotes “without” while “w/” denotes “with”. The best results for each series of SNNs are highlighted in bold font. $\uparrow$ (indicating "higher is better") and $\downarrow$ (indicating "lower is better") are used to denote performance metrics. Results highlighted with shading are ours. All results are averaged across $3$ random seeds.
Models PE Metric Metr-la (Short) Pems-bay (Short) Solar (Long) Electricity (Long) Avg.
@@
QKFormer w/o PE (Original) R2$\uparrow$ .$702$$501$$377$$240$$750$$702$$664$$655$$905$$751$$497$$375$$964$$963$$958$$955$685.685.685.685
@@
QKFormer w/ Conv-PE A R2$\uparrow$ .$717$$513$$376$$246$$767$$706$$681$$654$$920$$748$$512$$416$$970$$967$$963$$958$695.695.695.695
@@
5 Experiments
5.1 Datasets
To evaluate the RPE capabilities of the compared models, we conduct experiments on two sequential tasks: time-series forecasting and text classification. Following Lv et al. (2024b), we choose $4$ real-world datasets for time-series forecasting: Metr-la (Li et al., 2017b), Pems-bay (Li et al., 2017b), Electricity (Lai et al., 2018), Solar (Lai et al., 2018). For text classification, we follow Lv et al. (2024a) and conduct experiments on six benchmark datasets: Movie Reviews (Pang & Lee, 2005), SST-2 (Socher et al., 2013), SST-5, Subj, ChnSenti, and Waimai. Additionally, to demonstrate the versatility of our RPE method in image processing, we perform patch-based image classification experiments on one static dataset, CIFAR, and one neuromorphic dataset, CIFAR10-DVS (Li et al., 2017a). The details of these datasets, metrics, and training hyper-parameters are provided in Appendix B.
@@ 5.2 Time-Series Forecasting
Time-series forecasting is a suitable task for evaluating the effectiveness of our proposed RPE method because it involves capturing the temporal dependencies between data points over time. We follow the framework outlined by Lv et al. (2024b) to conduct time-series forecasting experiments using SNN models. Specifically, we use Spikformer (Zhou et al., 2023b) and the current state-of-the-art (SOTA) model, QKFormer (Zhou et al., 2024), as the backbone architectures, both of which incorporate spiking self-attention mechanisms. We modify the spiking self-attention mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR. Subsequently, we integrate our proposed Gray-PE and its extended forms into these SNN architectures.
For the baseline comparisons, we consider the following: (1) spiking absolute PE and CPG-PE (Lv et al., 2024a), (2) the PE used in the original architectures, referred to as Conv-PE, and (3) the integration of Binary-PE into the XNOR-based SSA. To isolate the effects of the XNOR modification, we also perform experiments on Spikformer-XNOR and QKFormer-XNOR without any positional encoding, as well as with Conv-PE. We present the performance of the compared SNN models with various positional encoding methods in Table 1. The key findings are as follows:
Table 2: Accuracy (%) on $6$ text classification benchmarks. The best results of SNNs and ANNs are formatted in bold font. We set the sentence length to $128$ for the English dataset and $32$ for the Chinese dataset. Experimental results are averaged across $5$ random seeds.
@@
(1) Relative PE generally outperforms absolute PE in time-series forecasting tasks. Based on the comparison between the Transformer with RPE and sinusoidal-PE, as well as Spikformer-XNOR with Conv-PE and Gray-PE, we observe that relative PE consistently outperforms absolute PE, particularly on long-term time-series datasets such as Solar and Electricity.
(2) The XNOR modification does not impact the performance of the original SNN models. The average performance of Spikformer without PE is nearly identical to that of Spikformer-XNOR without PE. This suggests that our XNOR modification of the SSA does not affect the performance of the original SNN models. Consequently, we can attribute the observed performance improvements solely to the introduced positional encoding techniques, excluding any impact from the XNOR operation itself.
(3) Gray-PE and its extended form, Log-PE, enable Spikformer and QKFormer to achieve the best performance among their variants. CPG-PE is a spiking version of absolute PE designed for SNNs. Spikformer and QKFormer, when equipped with our proposed Gray-PE and Log-PE, consistently outperform all other corresponding variants, including CPG-PE and Binary-PE. Furthermore, we were surprised to find that Binary-PE fails to capture the relative positional information. This leads us to reconsider the underlying issue: if a binary encoding method cannot represent distances between certain positions, it is unsuitable as an RPE solution for SNNs. In contrast, Gray Code and Log-PE provide approximate binary encoding approaches that “partially” achieve this goal.
(4) For long input sequences, Log-PE is more effective than Gray-PE in capturing relative positional information. The input sequence length for Metr-la and Pems-bay is $12$, whereas, for Solar and Electricity, it is $168$. As an approximated approach of RPE, Gray-PE can only represent the distance between two indexes differing by $2^{n}$. This limitation means that SNNs using Gray-PE may experience a performance drop when processing longer input sequences. Furthermore, on the long-sequence datasets Solar and Electricity, SNNs equipped with Log-PE consistently outperform those with Gray-PE across nearly all prediction length settings. For instance, on the Solar dataset with a prediction length of $48$, QKFormer with Log-PE significantly outperforms QKFormer with Gray-PE, showing an average improvement of $0.038$ in R2. This result indicates that Log-PE is more effective for processing long input sequences.
5.3 Text Classification
Text classification is another representative task for which RPE plays an important role. Therefore, we conduct experiments to evaluate the performance of SNNs with RPE in text classification tasks, comparing it against other PE techniques to highlight its advantages in capturing complex linguistic structures and contextual nuances. For this part, we strictly follow the experimental setup used in the text classification experiments in Lv et al. (2024a), and the results are presented in Table 2.
@@ 5.3 Text Classification
Table 3: Accuracy (%) on image classification benchmarks. Numbers with denote our implementation, representing results that differ from those reported in the original papers. The best and second-best results are highlighted in bold and underlined formats, respectively. “R” denotes relative PE, while “A” denotes absolute PE. The results with shading are ours. Results are averaged across $4$ random seeds.
Model PE CIFAR10 CIFAR10-DVS CIFAR100 Avg.
Spike Type Param (M) Accuracy Param (M) Accuracy Param (M) Accuracy
Vision-Transformer A $9.32$ $\bm{96.73}$ $9.36$ $\bm{81.02}$
Spikformer w/o PE $8.00$ $93.98$ $1.99$ $76.60$ $8.04$ $74.02$ $81.53$
Spikformer w/ Conv-PE (Original) A $9.32$   $\;\;94.80^{*}$ $2.57$   $\;\;78.10^{*}$ $9.36$   $\;\;77.04^{*}$ $83.31$
Spikformer w/ CPG-PE A $8.17$ $95.06$ $2.06$ $\underline{78.40}$ $8.20$ $77.82$ $83.76$
Spikformer-XNOR w/o PE $8.00$ $94.14$ $1.99$ $76.70$ $8.04$ $74.16$ $81.67$
Spikformer-XNOR w/ Conv-PE A $9.32$ $95.20$ $2.57$ $77.90$ $9.36$ $78.04$ $83.71$
Spikformer-XNOR w/ Gray-PE 1D R $8.00$ $\underline{95.46}$ $1.99$ $77.90$ $8.04$ $\underline{78.12}$ $\underline{83.83}$
Spikformer-XNOR w/ Gray-PE 2D R $8.00$ $\bf 95.66$ $1.99$ $\bf 78.70$ $8.04$ $\bf 78.45$ $\bf 84.27$
QKFormer w/o PE (Original) $6.74$ $\;\;\underline{96.32}^{*}$ $1.50$   $\;\;\bf 83.40^{*}$ $6.76$   $\;\;\bf 80.90^{*}$ $\bf 86.87$
QKFormer w/ CPG-PE A $7.01$ $96.30$ $1.58$ $82.00$ $7.04$ $80.52$ $86.27$
QKFormer-XNOR w/o PE $6.74$ $96.28$ $1.50$ $82.90$ $6.76$ $80.80$ $86.66$
QKFormer-XNOR w/ Gray-PE 1D R $6.74$ $96.22$ $1.50$ $82.20$ $6.76$ $80.48$ $86.30$
QKFormer-XNOR w/ Gray-PE 2D R $6.74$ $\bf 96.36$ $1.50$ $\underline{83.10}$ $6.76$ $\underline{80.82}$ $\underline{86.76}$

### v2

@@ 4.2 XNOR-Based Spiking Self-Attention
In the original Transformer [12], the attention map is computed via the dot product between the query and key matrices, $\mathbf{AttnMap}=\mathbf{Q}\cdot\mathbf{K}^{\operatorname{T}}$, which effectively captures similarity of $\mathbf{Q}$ and $\mathbf{K}$. As mentioned above, in order to capture the relative distances of spiking matrices while effectively measuring the similarity, we design the XNOR-based SSA. Unlike the dot-product operation, XNOR accounts for both spiking state ($1$) and the resting state (00).
Formally, we modify Equation 4 as follows:
$\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K}),$ (8)
@@
where $\lnot$ denotes the Not operation, $\oplus$ denotes the XOR operation, and $D$ represents the channel dimension. Note that every token in $\mathbf{Q}$ will perform XOR with every token in $\mathbf{K}$, so we sum over the channel dimension $D$ to get $\mathbf{AttnMap}\in\mathbb{N}_{0}^{T\times L\times L}$, shown in Figure 2 (a). The scale factor $\sigma$ in Equation 4 should be set to a smaller value or treated as a learnable parameter, ensuring that the firing rate of $\mathcal{SN}$ does not become excessively large. We will empirically demonstrate that this XNOR modification does not negatively impact the performance of the vanilla spiking self-attention.
4.3 Gray-PE
We propose that the Gray Code can serve as an approximate approach to relative positional encoding for spiking Transformers. This is supported by the following Theorem 1:
Theorem 1.
@@ 4.3 Gray-PE
(Proof in Appendix A) For two position indices differing by $2^{n}(n\geq 0)$, their Gray Code representations have a consistent Hamming distance. Specifically, $\forall$ position $i$, we have: $d_{H}(G(i),G(i+2^{n}))=\begin{cases}1&\text{if }n=0,\\ 2&\text{if }n\geq 1.\end{cases}$ (9)
As illustrated in Figure 2 (b), the Hamming distance $d_{H}(0,1)$ and $d_{H}(1,2)$ both equal $1$ because their relative distance is $1$, i.e., $02^{n},n=$ 0. Similarly, $d_{H}(0,2)=d_{H}(1,3)$, and $d_{H}(0,4)=d_{H}(1,5)$, as their relative distances are the power of $2$. That said, Gray Code ensures the consistency of relative distance representations for every $2^{n}$ ($n\geq 0$) relative distance.
@@
For implementation, we concatenate the Gray Code representations of each position index to both the query matrix $\mathbf{Q}\in\mathbb{N}_{0}^{T\times L\times D}$ and key matrix $\mathbf{K}\in\mathbb{N}_{0}^{T\times L\times D}$, leaving the remaining operations unchanged. We use concatenation instead of addition because $\mathbf{Q}$ and $\mathbf{K}$ are spike matrices, and addition would compromise their binary nature. Formally, the attention map $\mathbf{AttnMap}$ will be:
$\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot([\mathbf{Q}\parallel G(\bm{l})]\oplus[\mathbf{K}\parallel G(\bm{l})]),$ (10)
@@
where $G(\cdot)$ represents the function that converts integers into their binary Gray Code representations. The vector $\bm{l}$ denotes an array of position indexes, specifically $[0,1,2,\dots,L-1]$, where $L$ is the sequence length of $\mathbf{Q}$ and $\mathbf{K}$. $\parallel$ denotes concatenation on the channel dimension $D$.
Notably, the binary nature of Gray Code (comprising only 00 and $1$) aligns intrinsically with the spike-based computation paradigm, avoiding the need for floating-point operations that impose significant implementation overhead on neuromorphic hardware.
4.4 Log-PE
@@
Although Gray-PE can partially capture relative distances, it faces significant challenges when the input sequence is long or when the downstream task is highly sensitive to long-range dependencies. For instance, when $L\geq 10^{2}$, the distinguishable range of relative distances under Gray-PE becomes constrained by its power-of-two quantization mechanism. To mitigate this, we propose Log-PE that integrates logarithmic positional bias into spiking-based self-attention. Specifically, we simulate Equation 5 and follow ALiBi [13] to directly add a pre-assigned relative position map, denoted as $\mathbf{R}_{ij}$, to the attention map produced by SSA:
@@
$\mathbf{AttnMap}=\left(\sum_{i=0}^{D}\lnot(\mathbf{Q}\oplus\mathbf{K})\right)+\mathbf{R}_{i,j},\;\text{where}\;\;\mathbf{R}_{i,j}=\begin{bmatrix}R_{i,j}\end{bmatrix}=\begin{bmatrix}\lceil\log_{2}(\frac{L-1}{|i-j|+1})\rceil\end{bmatrix}.$ (11)
Here, $\lceil$.$\rceil$ . ⌉ denotes the round-up function, $L$ is the sequence length, and $i,j$ is position indices.
@@
Figure 2 (c) shows an illustration of Log-PE. Since the original $\mathbf{AttnMap}$ is a matrix composed of non-negative integers, we aim to ensure accurate relative distance consistency while preserving the effectiveness of spiking self-attention. Theoretically, if we set the $R_{i,j}$ as $\frac{L-1}{|i-j|+1}$, we could obtain a complete RPE for the spiking Transformers. However, we choose not to pursue this solution, because for long sequence lengths $L$, the large values of $\frac{L-1}{|i-j|+1}$ would catastrophically overshadow the original spiking attention activations (See Appendix B). Therefore, using the logarithmic form $R_{i,j}$ represents a compromise that balances the values between the spiking attention map and complete-RPE, while partially capturing relative position information.
@@ 4.5 Two-Dimensional Form for Image Patches
CNN-based SNN models, such as Spiking VGG [25] and SEW-ResNet [3], do not incorporate the concept of “positional encoding” in their spike representations. Vision Transformer [26] reformulated traditional image classification into a patch-based approach, dividing images into smaller patches. Unlike 1D positional encoding, which only considers the linear sequence of patches, 2D RPE accounts for both the horizontal and vertical positions of the patches in the image grid. This ensures that the model can recognize the relative positions along a single axis and the crucial interactions between patches across both dimensions. We show our 2D form in Figure 2 (d). In our implementation, we assign horizontal and vertical positions with independent dimensions to store the Gray Code. Formally, the attention map $\mathbf{AttnMap}$ is:
$\mathbf{AttnMap}=\sum_{i=0}^{D}\lnot\left([\mathbf{Q}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]\oplus[\mathbf{K}\parallel G(\mathbf{h})\parallel G(\mathbf{w})]\right).$ (12)
Here, $\mathbf{h}$ is the array of position indices, specifically $\mathbf{h}=[0,1,2,\dots,h-1]$, where $h$ denotes the maximum patch index along the height axis. Similarly, $\mathbf{w}$ is along the width axis. As for the 2D form of Log-PE, we can add $\mathbf{R^{h}}_{i,j}$ and $\mathbf{R^{w}}_{i,j}$ on $\mathbf{AttnMap}$, replacing the sequence length $L$ in Equation 11 with $h$ or $w$. However, in our pre-experiments, we found that spiking Transformers with Log-2D failed to converge due to the excessive magnitude. Therefore, we abandon the 2D form of Log-PE.
5 Experiments
5.1 Datasets
To evaluate the RPE capabilities of the compared models, we conduct experiments on two sequential tasks: time-series forecasting and text classification. Following [6], we choose $4$ real-world datasets for time-series forecasting: Metr-la [27], Pems-bay [27], Electricity [28], Solar [28]. For text classification, we follow [7] and conduct experiments on six benchmark datasets: Movie Reviews [29], SST-2 [30], SST-5, Subj, ChnSenti, and Waimai. Additionally, to demonstrate the versatility of our RPE method in image processing, we perform patch-based image classification experiments on two static datasets, CIFAR and Tiny-ImageNet, and one neuromorphic dataset, CIFAR10-DVS [2]. The details of these datasets, metrics, and training hyperparameters are provided in Appendix D.
Table 1: Experimental results of time-series forecasting on $4$ benchmarks with various prediction lengths $6,24,48,96$. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. “w/” denotes “with”. The best results for each series of spiking Transformers are highlighted in bold font. $\uparrow$ ($\downarrow$) indicates that the higher (lower) the better. Results highlighted with shading are ours. All results are averaged across $3$ random seeds.
Models PE Metric Metr-la ($L=12$) Pems-bay ($L=12$) Solar ($L=168$) Electricity ($L=168$) Avg.
@@ 5.1 Datasets
SDT-V1 w/ Conv-PE (Original) R2$\uparrow$ .$689$$517$$409$$253$$769$$700$$647$$630$$917$$819$$723$$655$$956$$952$$949$$950$721.721.721.721
@@
SDT-V1 w/ CPG-PE R2$\uparrow$ .$701$$525$$\bf.418$$257$$778$$\bf.716$$660$$\bf.656$$919$$\bf.820$$710$$644$$963$$960$$958$$952$727.727.727.727
[…]
R2$\uparrow$ .$\bf.714$$\bf.531$$415$$\bf.265$$\bf.784$$709$$\bf.672$$654$$\bf.921$$\bf.820$$\bf.730$$\bf.674$$\bf.972$$\bf.968$$\bf.963$$\bf.957$$\bf.734$
[…]
QKFormer w/ Conv-PE (Original) A R2$\uparrow$ .$717$$513$$376$$246$$767$$706$$681$$654$$920$$748$$512$$416$$970$$967$$963$$958$695.695.695.695
@@ 5.2 Time-Series Forecasting
We follow the SeqSNN [6] framework to conduct time-series forecasting experiments. Specifically, we take Spikformer [4], Spike-driven Transformer (SDT) V1 [5], and the current visual state-of-the-art (SOTA) model, QKFormer [20], as the backbone architectures. We modify the SSA mechanism as outlined in Section 4.2 to create two variants: Spikformer-XNOR and QKFormer-XNOR. SDT adopts a variant of SSA, which makes it only able to integrate Log-PE but not Gray-PE. We present the performance of the compared SNN models with various positional encoding methods in Table 1. The key findings are as follows:
(1) Directly applying RPE methods to spiking Transformers is ineffective. Specifically, Spikformer with rotary position embedding (RoPE) exhibits poor performance across all benchmarks. As discussed in Section 1, we argue that this limitation stems from the binary nature of spiking neurons during the computation of $\mathbf{Q}$ and $\mathbf{K}$, which makes it difficult to disentangle positional information from sparse spiking activations.
(2) The XNOR modification does not impact the performance of the original SNN models. The average performance of Spikformer with Conv-PE is nearly identical to that of Spikformer-XNOR with Conv-PE. This suggests that our XNOR modification of the SSA does not affect the performance of the original SNN models.
(3) Gray-PE and Log-PE, enable spiking Transformers to achieve the best performance among their variants. CPG-PE is a spiking version of absolute PE designed for SNNs. Spikformer and QKFormer, when equipped with our proposed Gray-PE and Log-PE, consistently outperform all other corresponding variants.
(4) For long input sequences, Log-PE is more effective than Gray-PE in capturing relative positional information. The input sequence length for Metr-la and Pems-bay is $12$, whereas for Solar and Electricity, it is $168$. On the long-sequence datasets Solar and Electricity, spiking Transformers equipped with Log-PE consistently outperform those with Gray-PE across nearly all prediction length settings. This result indicates that Log-PE is more effective for processing long input sequences.
5.3 Text Classification
We conduct experiments to assess the efficacy of spiking Transformers with Gray-PE and Log-PE in text classification tasks. By comparing them against alternative PE techniques, we demonstrate their superior ability to model complex linguistic structures and contextual dependencies. Our experimental setup strictly adheres to the methodology outlined in [7], and the results are shown in Table 2.
Table 2: Accuracy (%) on $6$ text classification benchmarks. Note that QKFormers fail to converge in the text classification task. Experimental results are averaged across $5$ random seeds.